# CoStruction: Conjoint radiance field optimization for urban scene reconStruction with limited image overlap

### v1

@@ Abstract
Reconstructing the surrounding surface geometry from recorded driving sequences poses a significant challenge due to the limited image overlap and complex topology of urban environments. SoTA neural implicit surface reconstruction methods often struggle in such setting, either failing due to small vision overlap or exhibiting suboptimal performance in accurately reconstructing both the surface and fine structures. To address these limitations, we introduce CoStruction, a novel hybrid implicit surface reconstruction method tailored for large driving sequences with limited camera overlap. CoStruction leverages cross-representation uncertainty estimation to filter out ambiguous geometry caused by limited observations. Our method performs joint optimization of both radiance fields in addition to guided sampling achieving accurate reconstruction of large areas along with fine structures in complex urban scenarios. Extensive evaluation on major driving datasets demonstrates the superiority of our approach in reconstructing large driving sequences with limited image overlap, outperforming concurrent SoTA methods.
@@
Figure 1: We introduce CoStruction, a novel hybrid implicit surface reconstruction method specifically designed for large-scale driving sequences with limited camera overlap. Extensive experiments on four major driving datasets demonstrate the superiority of CoStruction’s mesh (top) over previous state-of-the-art methods (bottom).
@@ 1 Introduction
Accurate 3D surface reconstruction of large urban scenes is essential for many challenging autonomous driving applications, such as scene relighting [25], simulation, and 3D object insertion [37]. However, achieving high-quality reconstructions in driving environments remains a significant challenge, where the difficulty primarily stems from two key factors:
@@
  • Geometric ambiguity: Limited camera overlap and the linear trajectory of vehicles introduce significant uncertainty in the reconstruction process.
@@
Neural Radiance Fields (NeRF) [21] have emerged as a powerful a powerful approach for 3D reconstruction. It represents 3D scenes as a volumetric density field using differentiable volume rendering equations enabling scene reconstruction from images and their associated camera poses. Although NeRF-based methods have demonstrated impressive results in capturing high frequency scene details  [1, 23], its under-constrained optimization problem leads to bad scene geometry, especially on planarregions [24, 5]. To adapt to large scene with complex structures, either 3D priors, such as LiDARs, or strong assumptions are introduced to constrain the origina l optimization problem [26, 8, 32, 31].
@@
Neural implicit surface methods [33, 34, 40], on the other hand, propose to replace the volumetric density field in NeRF with a Signed Distance Function (SDF). The SDF formulation represents the surface at its zero level sets and can be easily regularized using the eikonal constraint.Current neural SDF methods demonstrate high fidelity reconstruction quality in object-centric scenes with large overlap between the training images [17, 35, 41]. However, such formulation fail to reconstruct urban scenes captured by vehicle-mounted cameras due to limited observation overlap and linear sensor trajectory [7, 9].
To achieve high quality surface reconstruction for autonomous driving, recent methods explore a hybrid representation of the volumetric field and SDF field and have demonstrated promising reconstruction results [12, 9], but achieve suboptimal geometry reconstruction due to strong geometric assumptions about the scene topology or coarse and inaccurate initialization from volumetric field.
Following the idea of combining the strength of both representations, we propose to address the aforementioned challenges with CoStruction: a conjoint radiance field optimization solution using guided sampling via cross-representation uncertainty estimation. In particular, to address ambiguities in the scene geometry caused by limited camera overlap, we propose cross-representation uncertainty estimation to filter out invalid cues. Furthermore, to handle arbitrary object arrangements with large planar areas and fine structures, we propose to jointly optimize the two radiance fields via guided sampling for efficient rendering and accurate reconstruction. We evaluate our proposed solution on four public driving datasets: KITTI-360 [18], Pandaset [39], Waymo Open Dataset [29], and nuScenes [2], demonstrating that our method can accurately reconstructs complex urban geometry with limited image overlap.
@@
  • joint optimization of a volumetric and an SDF models with cross-supervision for efficient reconstruction of complex urban scene,
  • a novel guided sampling strategy leveraging cross-representation uncertainty estimation to filter out ambiguous geometric cues caused by limited overlap between training images,
  • adaptive masked Eikonal constrain to reduce over-regularization of SDF at early training stages.
@@ 2 Related work
Traditional surface reconstruction techniques, such as Multi-View Stereo (MVS) [22], have long been the cornerstone of 3D reconstruction tasks. However, their multi-step pipelines are prone to error accumulation, often resulting in incomplete or inaccurate reconstructions.
Neural implicit surface reconstruction methods have emerged as a promising solution to address these limitations. These methods primarily adapt volume rendering equations to achieve more accurate surface estimation [33, 40].Subsequent works have built upon this framework to reduce training times and enhance performance [34] or to improve the fidelity on complex surface reconstruction[17, 35] Notably, Neuralangelo [17] and Neurodin [35] achieve highly fidelity reconstruction of complex geometries in object-centric scenes where large observation overlaps are present. However, they face significant challenges when applied to autonomous driving scenarios [9, 7], where the camera trajectory is linear and image overlap is limited.
@@
To reconstruct surfaces in outdoor urban scenes for autonomous driving, recent methods often rely on 3D supervision [37, 27] or strong geometric priors [12, 31] to model large planar areas. StreetSurf [12] models different regions—close, far, and sky—using hierarchical space partitioning, 3D and 4D hash grids, and occupancy grids to facilitate efficient ray sampling. However, this approach is limited in applicability, as it imposes strong priors on the initialization of the Signed Distance Field (SDF) based on the vehicle’s ego poses, restricting its use to a narrower set of scenarios.
Recent approaches have also explored 3D Gaussian splatting for surface reconstruction [4, 11, 3]. These methods, however, often rely on heavy post-processing techniques, such as Poisson Surface Reconstruction [15], to extract the final surface, adding computational complexity. Gaussian Opacity Field (GoF) [42] offers a direct approach for mesh extraction by explicitly learning level sets, avoiding the need for post-processing. However, these methods may suffer from excessive memory usage when reconstructing large-scale driving scenes. To address this, StreetSurfGS [7] introduces a planar-based octree representation and segmented training, significantly reducing memory usage and making the method more suitable for large-scale scene reconstruction in autonomous driving.
@@
Newly developed hybrid approaches have sought to combine volumetric and Signed Distance Field (SDF) representations, aiming to leverage the strengths of both methods [30, 38, 9]. These methods demonstrate that integrating both volumetric and SDF representations can enhance both rendering efficiency and reconstruction accuracy. Specifically, Turki et al. [30] and Wang et al. [38] divide the scene into volumetric and surface regions based on a locally scaled SDF parameter, which is learned through an auxiliary head. They have shown promising results in achieving fast training and efficient rendering, especially for datasets with high overlap and object-centric scenes.
SCILLA [9], on the other hand, introduces a hybrid representation designed for large urban scene reconstruction. It utilizes a volumetric initialization of the signed distance field to progressively guide surface reconstruction, enabling the reconstruction of complex urban scenes with limited visual overlap. However, it still faces challenges in maintaining geometric accuracy, particularly in the reconstruction of fine details.
To address this limitation, we propose an improved approach that better combines the strengths of both representations through conjoint optimization, utilizing guided sampling via cross-representation uncertainty estimation.
@@ 3 Method
Figure 2: Overview of CoStruction.: we jointly train two implicit model: a volumetric representation $\mathcal{F}_{\Theta}^{v}$ and a SDF $\mathcal{F}_{\Theta}^{f}$ with a mutual guidance provided through our Guided Ray Sampling strategy. A colorized mesh of the scene $\mathcal{E}$ is periodically extracted from the SDF representation.
Given a collection of RGB images of an urban area with limited overlap and captured by sensors mounted on a moving vehicle, our goal is to resolve perspective geometry ambiguities introduced by partial observations of the scene and to reconstruct accurate surface with precise structural details.
3.1 CoStruction joint fields optimization
Overview.
We propose to jointly optimize two implicit representations to faithfully and efficiently reconstruct a complex urban scene: a volumetric radiance field $\mathcal{F}_{\Theta}^{v}$ and an SDF field $\mathcal{F}_{\Theta}^{f}$ with trainable weights $\Theta_{v}$ and $\Theta_{f}$, respectively. We also extract at regular intervals during the training process a colorized mesh $\mathcal{E}$ from $\mathcal{F}_{\Theta}^{f}$ using the marching cubes algorithm. An overview of our method is presented in Fig. 2.
We leverage a volumetric representation to achieve fast convergence and accurate fine-structure reconstruction. Indeed, as demonstrated in [9], volumetric radiance fields are both easier and faster to learn compared to SDFs. This is particularly crucial in outdoor urban scenarios, where the images used for reconstruction share small overlap between them, making it challenging to recover detailed scene geometry. To address this, we also jointly train an SDF representation to harness its advantages for precise surface reconstruction through the Eikonal constraint.
To train both representation efficiently, we introduce our main contribution: a Cross-Representation Guided Ray Sampling (GRS, Sec. 3.2). We propose leveraging cross-representation uncertainty estimation to filter out ambiguous information across representations. This uncertainty estimate is used to guide sampling for both representations, enabling precise surface reconstruction while preserving fine details. In the following paragraphs, we introduce major concept and notation related to our volumetric and SDF representation.
@@ 3.1 CoStruction joint fields optimization
A volumetric radiance field is a continuous function $\mathcal{F}_{\theta}^{v}$ mapping a position and direction pair$(\mathbf{x},\mathbf{u)}\in\mathbb{R}^{3}\times\mathbb{S}^{2}$ to a volume density $\sigma\in\mathbb{R}^{+}$ and a color value $\mathbf{c}\in[0,1]^{3}$. Mildenhall et al. [21] model this function with a multi layer perceptron (MLP) whose weights $\theta$ are optimized to reconstruct a 3D scene given a set of posed images during training.
To render an image and depth from NeRF, volume rendering is applied to alpha-composite the color for each ray, yielding the final pixel color $\hat{C}_{v}(r)\in\mathbb{R}^{3}$ and the depth value $\hat{D}_{v}(r)\in\mathbb{R}$ to be:
$\hat{C}_{v}(r)=\sum_{i=1}^{N}w_{i}c_{i},\ \ w_{i}=T_{i}\alpha_{i},$ (1)
$\hat{D}_{v}(r)=\sum_{i=1}^{N}w_{i}z_{i},\ \ w_{i}=T_{i}\alpha_{i},$ (2)
where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the accumulated transmittance, $\alpha_{i}\in\mathbb{R}$ the blending factor and $z_{i}$ is the distance of the sample to the camera center. Here, $\alpha_{i}$ is computed from the predicted density: $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$, with $\delta_{i}\in\mathbb{R}$ being the distance between the samples along the ray.
@@
To constraint the optimization for better surface reconstruction, SDF-based approaches represent the scene surfaces as the zero level set of a signed distance function [33, 40]: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. To enable volume rendering, the signed distance function $f$ is then transformed into the density $\sigma$ using sigmoid-shape mapping for alpha compositing..
NeuS [33] adopted a new formulation of the blending factor $\alpha$:
$\alpha_{i}=\text{max}\left(\frac{\Phi_{s}(f(p_{i}))-\ \Phi_{i}(f(p_{i+1}))}{\Phi_{s}(f(p_{i}))},0\right),$ (3)
where $f(p_{i})$ and $f(p_{i+1})$ are signed distance values at section points centered on $x_{i}$, $\Phi_{s}(x)$ the sigmoid function. Such formulation is unbiased in depth estimation, which enables high fidelity surface reconstruction [17].
3.2 Guided ray sampling
Overview.
Because both of our representation use volumetric rendering to generate images, we decide to cross-supervise each model within the ray sampling stage with our Guided Ray Sampling (GRS) strategy. GRS can be divided into four steps: 1) ray generation, 2) per-ray uncertainty computation, 3) ray classification, and 4) ray sampling. Similar to RobustNeRF [28], we use a robust kernel to compute a per-ray uncertainty value. These values are used to compute uncertainty masks $\mathcal{M}_{d,c}$ to filter rays for the subsequent ray sampling step. Below, we describe more in details how uncertainty values are computed and we introduce the GRS applied to the volumetric model and the one used with the SDF representation.
@@ 3.2 Guided ray sampling
Due to partial observations in driving setting, each representation has limitations in capturing certain geometric properties to due limited vision overlap. Therefore, detecting uncertain predictions is crucial for determining which representation to trust.To address this, we leverage depth and photometric uncertainty across predictions from $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$ to identify and filter out noisy predictions.
One intuitive approach to estimate the uncertainty is to compare the expected ray $r$ termination (depth) according to $\mathcal{F}_{\Theta}^{v}$, $\hat{D}_{v}(r)$, and the distance to the first intersection of $r$ with the mesh $\mathcal{E}$, $\hat{D}_{\mathcal{E}}(r)$:
$\mathcal{M}_{d}(r)=\begin{cases}1,&\text{if }|1-\frac{\hat{D}_{\mathcal{E}}(r)}{\hat{D}_{v}(r)}|<\tau_{d}\\ 0,&\text{otherwise}\end{cases}$ (4)
where $\tau_{d}$ is predefined threshold. We introduce a second uncertainty mask, based on a photometric criteria, using the color of the mesh $\mathcal{E}$ associated to ray $r$, $\hat{C}_{\mathcal{E}}(r)$, defined as:
$\mathcal{M}_{c}(r)=\begin{cases}1,&\text{if }|\hat{C}_{\mathcal{E}}(r)-C_{gt}(r)|<\tau_{c}\\ 0,&\text{otherwise}\end{cases}$ (5)
where $C_{gt}(r)$ is the ground truth RGB values associated to ray $r$ and $\tau_{c}$ is another threshold parameter. The uncertainty mask $\mathcal{M}_{d}$ provide a measure of the joint convergence of $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$ while $\mathcal{M}_{c}$ indicate part of the scene where the SDF representation $\mathcal{F}_{\Theta}^{f}$ is close to the GT geometry.
GRS for volumetric representation.
To improve the volumetric reconstruction on planar surfaces, we guide the volumetric process according to $\mathcal{M}_{c}$. The mask $\mathcal{M}_{c}$ provide a simple yet effective indicator to divide the scenes according to the photometric accuracy of SDF field. Empirically the mask will indicate region with high frequency details and fine structures.
Given a batch of rays $r\in\mathcal{R}_{v}$ emitted from the camera and $\left[t_{n},t_{f}\right]$ be the original sampling interval of a ray $r$. To guide the sampling of the volumetric representation with the SDF, we adjust the sampling bounds according to $\hat{D}_{\mathcal{E}}(r)$:

### v2

# J-NeuS: Joint field optimization for Neural Surface reconstruction in urban scenes with limited image overlap

@@ Abstract
Reconstructing the surrounding surface geometry from recorded driving sequences poses a significant challenge due to the limited image overlap and complex topology of urban environments. SoTA neural implicit surface reconstruction methods often struggle in such setting, either failing due to small vision overlap or exhibiting suboptimal performance in accurately reconstructing both the surface and fine structures. To address these limitations, we introduce J-NeuS, a novel hybrid implicit surface reconstruction method for large driving sequences with outward facing camera poses. J-NeuS leverages cross-representation uncertainty estimation to tackle ambiguous geometry caused by limited observations. Our method performs joint optimization of two radiance fields in addition to guided sampling achieving accurate reconstruction of large areas along with fine structures in complex urban scenarios. Extensive evaluation on major driving datasets demonstrates the superiority of our approach in reconstructing large driving sequences with limited image overlap, outperforming concurrent SoTA methods.
@@
Figure 1: We introduce J-NeuS, a novel hybrid implicit surface reconstruction method specifically designed for large-scale driving sequences with limited camera overlap. Extensive experiments on four major driving datasets demonstrate the superiority of J-NeuS’s mesh (top, form left to right: mesh normals, textured mesh and shaded mesh) over previous state-of-the-art methods (bottom).
@@ 1 Introduction
Accurate 3D surface reconstruction of large urban scenes is essential for many challenging autonomous driving applications, such as scene relighting [25], sensor simulation [40], and 3D object insertion [37]. However, achieving high-quality reconstructions in driving environments remains a significant challenge, where the difficulty primarily stems from two key factors:
@@
  • Geometric ambiguity from outward facing sensors: Limited camera overlap and the linear trajectory of vehicles introduce significant uncertainty in the reconstruction process.
@@
Neural Radiance Fields (NeRF) [21] have emerged as a powerful approach for 3D scene reconstruction in such settings. Leveraging differentiable volume rendering, NeRF enables accurate reconstruction from images and their associated camera poses, effectively avoiding the error accumulation typically seen in traditional multi-view stereo pipelines. Although NeRF-based methods have demonstrated impressive results in capturing high frequency scene details  [1, 23], its under-constrained optimization problem leads to bad scene geometry, especially when vision cues are limited [24, 5]. To adapt to large scenes with complex structures, either 3D priors, such as LiDAR pointcloud, or strong assumptions are introduced to constrain the original optimization problem [26, 8, 32, 31].
@@
Neural implicit surface methods [33, 34, 41], overcome the limitation of NeRF by replacing the volumetric density field with a Signed Distance Function (SDF). The SDF formulation represents the surface at its zero level sets and is regularized using the Eikonal constraint. Current neural SDF methods demonstrate high fidelity reconstruction quality in object-centric scenes with large texture-less surfaces [17, 42]. However prevalent SDF methods like Neuralangelo [17] work better on landmark-centric scenes with large image overlaps but fail to reconstruct urban scenes captured by vehicle-mounted cameras due to limited observation overlap [7, 9]. Moreover, SDF-based methods often struggle to preserve fine structural details due to biased depth estimation and over-regularization of geometric constraints [35, 44], an issue that is critical for downstream autonomous driving applications.
To achieve high-quality surface reconstruction for autonomous driving – capturing both fine structures and large surfaces – recent approaches integrate volumetric and SDF representations. These methods partition scenes into distinct regions and apply specialized reconstruction [12] or sampling strategies [30, 38] tailored to each region’s unique characteristics. Additionally, other methods [9] demonstrate that SDF representations can benefit from volumetric initialization, enabling faster convergence and improved geometric fidelity. However, these solutions often result in suboptimal geometry quality due to overly strong geometric assumptions about the scene or reliance on coarse initialization that introduce noise from volumetric prediction.
In order to accurately combine the strengths of both volumetric and SDF representation, we propose an uncertainty estimation framework that identifies a noisy predictions arising from geometric ambiguity. Specifically, we estimate two types of uncertainty – photometric uncertainty and geometric uncertainty – to jointly train the volumetric and SDF models. These uncertainty measures guide the sampling process, ensuring that each representation is deployed where it performs better while mitigating over-regularization in the SDF to facilitate fine structure learning. Our approach effectively handles the complex geometry of urban environments, enabling efficient rendering and precise surface reconstruction. We evaluate our proposed solution on four public driving datasets: KITTI-360 [18], Pandaset [39], Waymo Open Dataset [29], and nuScenes [2], demonstrating robust reconstructions of intricate urban geometry with limited image overlap.
@@
  • Joint optimization of NeRF and SDF: We propose a framework that fuses volumetric (NeRF) and surface (SDF) representations so each excels in regions where it is better.
  • Guided Ray Sampling: We introduce a novel sampling strategy that leverages cross-representation uncertainty to tackle ambiguous geometric cues, enabling faster and more accurate surface reconstruction.
  • Adaptive Relaxation on geometry regularization: We dynamically relax the Eikonal constraint and monocular cues in uncertain regions to avoid over-smoothing, ensuring complete fine-structure reconstruction.
@@ 2 Related work
Traditional surface reconstruction techniques, such as Multi-View Stereo (MVS) [22], have long been the cornerstone of 3D reconstruction tasks. However, their multi-step pipelines are prone to error accumulation, often resulting in incomplete or inaccurate reconstructions. In contrast, neural implicit surface reconstruction approaches adapt volume rendering for more accurate surface estimation [33, 41], reducing the need for manual post-processing. Subsequent advances target faster training [34] or to and more complex geometry [17, 35]. Notably, Neuralangelo [17] and Neurodin [35] achieve highly fidelity reconstruction of complex geometries in object-centric scenes where large observation overlaps are present. However, they face significant challenges when applied to autonomous driving scenarios [9, 7], where the camera trajectory is linear and image overlap is limited.
@@
To enhance surface reconstruction in autonomous driving scenes with limited image overlap, recent methods typically rely on 3D supervision [37, 27], strong geometric priors [12, 31], or monocular supervision [12, 9]. While these approaches effectively model large planar areas, they often struggle to capture fine details due to strong geometric assumptions [31], inconsistencies in monocular predictions [12] or over-regularization on intricate features [9].
Recent approaches have also explored 3D Gaussian splatting for surface reconstruction [4, 11, 3]. These methods, however, often rely on heavy post-processing techniques, such as Poisson Surface Reconstruction [15], to extract the final surface, adding computational complexity. Gaussian Opacity Field (GoF) [43] offers a direct approach for mesh extraction by explicitly learning level sets, avoiding the need for post-processing. However, these methods may suffer from excessive memory usage when reconstructing large-scale driving scenes. To address this, StreetSurfGS [7] introduces a planar-based octree representation and segmented training, significantly reducing memory usage and making the method more suitable for large-scale scene reconstruction in autonomous driving. Nonetheless, it still requires TSDF fusion and additional mesh cleaning to achieve acceptable mesh quality.
@@
To overcome the limitations of SDF methods in capturing fine structures, recent approaches integrate volumetric and SDF representations by dividing scenes into distinct regions and applying specialized reconstruction or sampling strategies. Turki et al. [30] and Wang et al. [38] segment scenes based on the scaling factor used to convert SDF into density, and employ distinct sampling techniques for volumetric and surface regions to enable real-time, high-quality rendering. StreetSurf [12] models different regions – close, far, and sky – using different hierarchical space partitioning (e.g., 3D/4D hash grids, occupancy grids). This improves performance in urban scenes but imposes strong priors tied to vehicle ego poses, limiting generalizability. Meanwhile, ViiNeuS [9] initializes SDF sampling with volumetric density before gradually transitioning to a pure SDF representation, accelerating convergence and achieving state-of-the-art results on autonomous driving benchmarks. However, this approach risks introducing artifacts from the volumetric representation that can adversely affect the final SDF reconstruction.
Building on these insights, we propose a more adaptive divide-and-conquer technique that dynamically partitions the scene based on uncertainty across both representations. By applying tailored sampling and regularization strategies for different region, our method preserves fine details while effectively handling the large planar areas characteristic of autonomous driving environments.
@@ 3 Method
Figure 2: Overview of J-NeuS: we jointly train two implicit model: a volumetric representation $\mathcal{F}_{\Theta}^{v}$ and a SDF $\mathcal{F}_{\Theta}^{f}$ with a mutual guidance provided through our Guided Ray Sampling strategy. A colorized mesh of the scene $\mathcal{E}$ is periodically extracted from the SDF representation.
Given a collection of RGB images captured from a moving vehicle in an urban area with limited overlap, our goal is to resolve perspective geometry ambiguities introduced by partial scene observations and to accurately reconstruct surfaces with precise structural details. To achieve this, we propose selectively fusing volumetric and SDF representations, with uncertainty guiding the division of labor between the two representations. Specifically, we use two implicit representations: a volumetric radiance field $\mathcal{F}_{\Theta}^{v}$ and an SDF field $\mathcal{F}_{\Theta}^{f}$ with trainable weights $\Theta_{v}$ and $\Theta_{f}$, respectively. We also extract at regular intervals during training a colorized mesh $\mathcal{E}$ from $\mathcal{F}_{\Theta}^{f}$ using the marching cubes algorithm. An overview of our method is presented in Fig. 2
To selectively fuse NeRF and SDF representations during training, we first introduce photometric and geometric uncertainty estimation across the two representations $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$ (Section. 3.1). We then jointly optimize both representations with Guided Ray Sampling based on uncertainty estimation to leverage the strengths of both representations (Section. 3.2). Additionally, to ensure the preservation of fine structures, we relax the Eikonal constraint and monocular supervision in uncertain regions to avoid over-regularization of the SDF field (Section. 3.3).
3.1 Cross Representation Uncertainty Estimation
Due to the inherent partial observations in autonomous driving scenarios, both volumetric and SDF representations exhibit epistemic uncertainty in regions where visual cues are sparse. These regions are characterized by high variance in RGB and depth predictions, or deviations from the ground truth. To fully leverage the strengths of both representations, it is crucial to identify areas with high uncertainty and adaptively apply tailored sampling and regularization strategies. In the following sections, we first introduce the key concepts and notation related to our volumetric and SDF representations. We then describe how uncertainty is estimated to selectively fuse both representations effectively.
@@ 3.1 Cross Representation Uncertainty Estimation
[…]
To render an image and depth, volume rendering is applied to alpha-composite the color for each ray, yielding the final pixel color $\hat{C}_{v}(\mathbf{r})\in\mathbb{R}^{3}$ and the depth value $\hat{D}_{v}(\mathbf{r})\in\mathbb{R}^{+}$ to be:
$\hat{C}_{v}(\mathbf{r})=\sum_{i=1}^{N}w_{i}c_{i},\ \ w_{i}=T_{i}\alpha_{i},$ (1)
$\hat{D}_{v}(\mathbf{r})=\sum_{i=1}^{N}w_{i}z_{i},\ \ w_{i}=T_{i}\alpha_{i},$ (2)
where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the accumulated transmittance, $\alpha_{i}\in\mathbb{R}$ the blending factor and $z_{i}\in\mathbb{R}^{+}$ is the distance of the sample to the camera center. Here, $\alpha_{i}$ is computed from the predicted density: $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$, with $\delta_{i}\in\mathbb{R}^{+}$ being the distance between samples along the ray.
@@
While the volumetric field predicts a pointwise 3D density $\sigma(x)$ at each location $x$, the SDF instead predicts a signed distance function $f(x)$ and converts it to density using a logistic function $\phi_{s}(f)$ with a global scale parameter $s$, effectively confining the density to a narrow band of width $O(1/s)$ around the surface [38]. To enable volume rendering, the signed distance function $f$ is then transformed into the density $\sigma$ using sigmoid-shape mapping for alpha compositing. NeuS [33] adopted a new formulation of the blending factor:
$\alpha_{i}=\text{max}\left(\frac{\Phi_{s}(f(p_{i}))-\ \Phi_{s}(f(p_{i+1}))}{\Phi_{s}(f(p_{i}))},0\right),$ (3)
where $f(p_{i})$ and $f(p_{i+1})$ are signed distance values at section points centered on $x_{i}$, $\Phi_{s}(x)$ the sigmoid function.
@@
When reconstructing urban scenes, the direct 3D point-based density prediction of the volumetric field enables rapid fitting of high-frequency geometry (e.g., poles), but may also introduce spurious density (floaters) on large planar surfaces [33]. In contrast, the more constrained density prediction of the SDF promotes surface continuity, but may over-smooth fine details when using a uniform or overly small scale parameter $s$. To harness the complementary strengths of both representations, it is essential to identify and suppress their unreliable predictions. To this end, we introduce two complementary uncertainty estimates: geometric uncertainty ($\mu_{d}$) and photometric uncertainty ($\mu_{c}$).
The first type of uncertainty focuses on identifying regions with ambiguous geometric information: areas where visual cues are insufficient for geometry reconstruction (e.g., textureless surfaces or partially observed complex structures). Ideally, one would measure epistemic uncertainty by comparing predictions with dense 3D ground truth. However, in autonomous driving scenarios, such ground truth is often unavailable. Therefore, we propose to measure uncertainty by evaluating the consistency between the two representations: Given a ray $r$, we introduce the geometry uncertainty based on the rendered depth from $\mathcal{F}_{\Theta}^{v}$, and the distance to the first intersection with the mesh $\mathcal{E}$, $\hat{D}_{\mathcal{E}}(\mathbf{r})$:
$\mathcal{\mu}_{d}(\mathbf{r})=|1-\frac{\hat{D}_{\mathcal{E}}(\mathbf{r})}{\hat{D}_{v}(\mathbf{r})}|.$ (4)
High values of $\mu_{d}$ indicate inconsistency between the volumetric and SDF models, suggesting that at least one representation is uncertain. In such cases, we preferentially rely on the volumetric field for two key reasons. First, during initialization, it captures fine structures more completely and achieves faster convergence [9]. Second, in later stages, probability sampling 3.4 of $\mathcal{F}_{\Theta}^{f}$ can effectively filter out noise introduced by the volumetric field $\mathcal{F}_{\Theta}^{v}$ while preserving the fine details that $\mathcal{F}_{\Theta}^{v}$ initially localized.
Secondly, we introduce a photometric-based uncertainty to evaluate the predictions of the SDF field. We hypothesize that if the learned SDF field is accurate, a single-sample rendering at the mesh depth should yield a precise photometric rendering. Casting a ray $\mathbf{r}=\mathbf{o}+t\mathbf{u}$ from the camera center $\mathbf{o}$ through the pixel along direction $\mathbf{u}$, we define the photometric uncertainty indicator as:
$\mathcal{\mu}_{c}(\mathbf{r})=|\hat{C}_{\mathcal{E}}(\mathbf{r})-\bar{C}(\mathbf{r})|,$ (5)
where $\bar{C}$ indicates rgb ground truth, and $\hat{C}_{\mathcal{E}}(\mathbf{r})$ represents the rgb value of $\mathcal{F}_{\Theta}^{f}$ at the point $\mathbf{o}+\hat{D}_{\mathcal{E}}(\mathbf{r})\mathbf{u}$, i.e. the point on the Mesh $\mathcal{E}$ of the given ray. $\mathcal{\mu}_{c}(\mathbf{r})$ is used to divide certain and uncertain regions in the SDF representation (see Section. 3.3). We use it to guide the sampling strategy of the $\mathcal{F}_{\theta}^{v}$ for certain regions, and adaptively relax geometric regularization in uncertain regions to prevent over-smoothing. While $\mu_{c}$ does not guarantee accurate depth alignment, it is sufficient to quickly populate large planar regions without requiring complete rendering of $\mathcal{F}_{\theta}^{v}$. We provide a visualization of $\mu_{c}$ and $\mu_{d}$ in Figure 3.
$\mu_{d}$ $\mu_{c}$
Epoch 01 Refer to caption Refer to caption
Epoch 11 Refer to caption Refer to caption
GT. RGB Refer to caption
Figure 3: Visualization of $\mu_{d}$ and $\mu_{c}$, maximum clamped at 0.3.
3.2 Guided Ray Sampling (GRS)
To selectively fuse the volumetric and SDF representation during training, we propose to jointly train $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$ via Guided Ray Sampling (GRS) based on previously introduced uncertainty $\mathcal{\mu}_{c}$ and $\mathcal{\mu}_{d}$. Specifically, for the volumetric representation, we use $\mathcal{\mu}_{c}$ to guide the sampling toward the surface of the mesh $\mathcal{E}$ where SDF field $\mathcal{F}_{\Theta}^{f}$ is confident. For the SDF representation, we focus the sampling to the estimated surface of the best-suited representation, using geometric uncertainty $\mathcal{\mu}_{d}$.
Volumetric-GRS.
We consider a batch of rays $r\in\mathcal{R}_{v}$ with for each ray $r$ the corresponding original sampling interval $\left[t_{n},t_{f}\right]$. To improve volumetric reconstruction on planar surfaces, we adjust the sampling bounds to focus more on the estimated depth of the Mesh $D_{\mathcal{E}}(\mathbf{r})$ if the $\mathcal{\mu}_{c}(\mathbf{r})$ is below a certain threshold $\tau_{c}$: